Nothere:

A 10-ounce container of chocolate chips can be used to make 4 desserts. Each dessert has the same number of chocolate chips. Which equation shows how to find the number of ounces in each dessert? 10÷4=104 ounces 4÷10=104 ounces 10÷4=410 ounce 4÷10=410 ounce

jhonyy9:

using the rule of 3 simple you can write 10 ounce container of chocolate chips use to make 4 desserts x ounce container ... to make 1 dessert cross multiplie and calcule the x value x = 10*1 /4 = 10 / 4 hope helped understandably easy